This is the final question I have left and I just can't seem to figure it out.
Here it is.
Four years from now, Kateyln will be as old as Adrian is now. Seven years ago, the sum of their ages was 22. Write a system of equations and solve is using elimination to determine how old the girls are now.

Answer by ankor@dixie-net.com(22740)

let k = Kate's present age
let a = Adrian present age
:
Write an equation for each statement
:
Four years from now, Kateyln will be as old as Adrian is now.
k + 4 = a
this form for elimination
a - k = 4
:
Seven years ago, the sum of their ages was 22.
(a-7) + k-7) = 22
a + k = 22 + 14
a + k = 36
:
a - k = 4
a + k = 36
--------------Addition eliminates k, find a
2a = 40
a = 20 yrs is Adrian's age
:
You can find Kate's age, check your solutions in both statements



Write a system of equations and solve is using elimination to determine how old the girls are now.